OpenStudy (zyberg):

Could anyone check my solution? On a resistance usually is written a maximum power that should not be crossed over. On a resistance is written: 50 k Ohms, 2 W. What is the maximum electrical current that is possible to reach with this resistor?

OpenStudy (zyberg):

Since P = I^2 * R; I = sq root (P/R); I = sq root (50 000 Ohms/2 W) = sq root (25 000) ~ 158.11 Amperes.

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (zyberg):

Woah! I was very unsure because of the number being somewhat random. Thank you!

ganeshie8 (ganeshie8):

No wait, you flipped the ratio

ganeshie8 (ganeshie8):

`Since P = I^2 * R; I = sq root (P/R); I = sq root (50 000 Ohms/2 W) ` Shouldn't that be sq root (2 W / 50 000 Ohms) ?

OpenStudy (zyberg):

Oh. Seems that my instincts to divide bigger number by a lesser got ahead of me. Thank you for spotting this :)

ganeshie8 (ganeshie8):

Np, so the max current allowed seems to be 6.3 milli amperes